Cycle Business manufactures and sells road and mountain bikes through a network of retail outlets in western Canada. Below is a partial list of expense items incurred in the most recent month (November), when 1,160 bicycles were manufactured, shipped, and sold. There was no beginning or ending work in process or finished goods inventory in November Item Units produced and sold Sales October November 1,060 1,160 Leather used for the bicycle seats Production managers salary Life insurance for the company president Electricity used in the production facilities Sales commissions Internet advertising Employee benefits for the production workerst Property taxes on the production facilities Shipping costs Salary of the chief financial officer $1,060,000 $1,160,000 41.760 6,800 280 2,500 51,920 1,300 41,760 1,800 51,920 11,600 38,160 6,800 280 2,300 46,600 1,300 38,160 1,800 46,600 11,600 Each month, regardless of how much electricity is used, Cycle Business pays a $180 base charge to the utilities company Employee benefits total 20% of the wages paid to production workers, who on average earn $20 per hour. Each bicycle requires 9 hours of direct labour. Required 1. With respect to the partial list of November expenses a. Use the Type column to label each item as a Manufacturing, Administrative, or Selling cost. b. For those items you labeled as a Manufacturing cost, i. Use the Behaviour column to label each item as variable or fixed manufacturing costs ii Use the Direct/lndirect column to label each item as a direct or indirect cost Item Type Behaviour Direct/lndirect Leather used for the bicycle seats Production managers salary Life insurance for the company president Electricity used in the production facilities Sales commissions Internet advertising Employee benefits for the production workers Property taxes on the production facilities Shipping costs Salary of the chief financial officer rect e/fixed Selling Selling Manufacturing anufacturing
2. Assume that 1,360 bicycles will be manufactured and sold in December. For the items you classified as manufacturing costs in requirement 1, estimate the cost for December. Assume that there will be no change in unit costs for any direct materials, hourly wages will remain the same, and employee benefits will continue at 20% of wages. Cost
Answer #1

Cost of 1360 bicycles is as follows :

Leather used for the bicycle seats = 1360*36# = $48,960

#(38160/1060) or (41760/1160)

Production managers salary = $6,800

Electricity used in the production facilities = 180 + (1360*2#) = $2,900

#(2300-180)/1060 or (2500-180)/1160

Employee benefit for production workers = 1360*20*9*20% = $48,960

Property tax on production facilities = $1,800

So, the total cost is $109,420